### v3

@@ 4.5 Ablation Study
Effect of DAQ and ICL. We implement two ablated models to study the efficacy of the proposed DAQ and ICL designs. Tab. 4 presents the results on RepCountA [14] using the TSN backbone. In ablation (a), we substitute the proposed DAQ module with a static action query, where the action queries $\mathbf{Q}^{\text{act}}$ are also learnable variables, instead of using the action queries filtered out by the encoder and query selection modules. In ablation (b), we eliminate the ICL design among the action queries. The results demonstrate that both DAQ and ICL contribute to performance improvement, particularly in terms of counting medium and long actions.
Confidence threshold $\alpha$. We report the impact of different confidence thresholds for classifying “repetitive actions” on the final counts in Fig. 6. We illustrate the MAE (left) and OBO (right) curves for our method (green curve) and the TransRAC [14] approach (blue curve) across various confidence thresholds $\alpha$. The metrics for TransRAC are derived by binarizing its output density map and summing the results to obtain the final count. The results indicate that setting the threshold within the range of $0.2$ to $0.4$ yields similar performance for our method. Our method consistently outperforms TransRAC [14] (blue curve) by a considerable margin.
Figure 6: Results of different confidence thresholds of our method and TransRAC [14]. We depict the MAE (left) and OBO (right) curves of our method (green curve) and the TransRAC (blue curve) approach concerning different confidence thresholds. The metrics of TransRAC are obtained by binarizing the density map of TransRAC output and then summing to obtain the final count.
@@ 5 Conclusion
Data Availability Statements. The RepCountA [14] and UCFRep [47] datasets that support the findings of this study are publicly available on GitHub at https://github.com/SvipRepetitionCounting/TransRAC and https://github.com/Xiaodomgdomg/Deep-Temporal-Repetition-Counting, respectively.

### v4

@@ 4.5 Ablation Study
Effect of DAQ and ICL. We implement two ablated models to study the efficacy of the proposed DAQ and ICL designs. Tab.˜4 presents the results on RepCountA [hu2022transrac] using the TSN backbone. In ablation (a), we substitute the proposed DAQ module with a static action query, where the action queries $\mathbf{Q}^{\text{act}}$ are also learnable variables, instead of using the action queries filtered out by the encoder and query selection modules. In ablation (b), we eliminate the ICL design among the action queries. The results demonstrate that both DAQ and ICL contribute to performance improvement, particularly in terms of counting medium and long actions.
Confidence threshold $\alpha$. We report the impact of different confidence thresholds, which range from 0 to 1 for classifying “repetitive actions” on the final counts in Fig.˜6. During inference, queries with prediction confidence scores below the threshold are classified as “others”, while the rest are considered “repetitive actions”. We illustrate the MAE (left) and OBO (right) curves for our method (green curve) and the TransRAC [hu2022transrac] approach (blue curve) across various confidence thresholds $\alpha$. The metrics for TransRAC are derived by binarizing its output density map and summing the results to obtain the final count.
As the figure illustrates, our model performance drops significantly when the threshold exceeds 0.5, e.g. OBO becomes nearly zero, since most queries are mistakenly filtered out. This aligns with our expectations. Conversely, when the threshold is set between 0.3 and 0.5, the model yields consistently strong performance. However, setting the threshold too low also leads to performance degradation, as nearly all queries are then classified as “repetitive actions”, including irrelevant or noisy ones. This highlights the importance of selecting an appropriate threshold range to achieve optimal performance, as the results indicate that setting the threshold within the range of $0.2$ to $0.4$ yields consistently strong performance for our method. Besides, our method consistently outperforms TransRAC [hu2022transrac] (blue curve) by a considerable margin.
Figure 6: Results of different confidence thresholds of our method and TransRAC [hu2022transrac]. We depict the MAE (left) and OBO (right) curves of our method (green curve) and the TransRAC (blue curve) approach concerning different confidence thresholds. The metrics of TransRAC are obtained by binarizing the density map of TransRAC output and then summing to obtain the final count.
@@ 5 Conclusion
Data Availability Statements. The RepCountA [hu2022transrac] and UCFRep [zhang2020context] datasets that support the findings of this study are publicly available on GitHub at https://github.com/SvipRepetitionCounting/TransRAC and https://github.com/Xiaodomgdomg/Deep-Temporal-Repetition-Counting, respectively.